An astronomical unit is the average distance Of The earth to the sun, approximately 150,000,000 km. The speed of light is about 3 X 10^8m/s. Express the _peed of light in terms of astronomical units per minute
1
Expert's answer
2020-12-28T08:43:41-0500

Find how much time it will take to pass 150 000 000 000 m (1 AU) at the speed of light:


"t=\frac{150\cdot10^9}{3\cdot10^8}=500\text{ s}."

Thus, the speed of light is 1 AU per 500 seconds. In other words, it is 1 AU in 500/60=8.33 min. In other words, it is 1 AU/8.33 min=0.12 AU/min.